# Aligning Human and LLM Judgments: Insights from EvalAssist on Task-Specific Evaluations and AI-assisted Assessment Strategy Preferences

### v1

@@ 5.1. Participant Demographics
We recruited 15 participants at ABC company who had prior experience with model evaluation. The job titles are listed in Table 3. Their prior experience included subjective and manual evaluation, use case specific evaluation, evaluation framework and metrics.
@@
5.2. RQ1: What do practitioners prioritize in evaluation criteria when using LLMs as judges, and how do their priorities differ?
@@
The criteria that users created varied widely across tasks, not only in terms of specificity, but also in the presence of instructions or rules, the number of items in the scale (for the direct assessment condition), the inclusion of exclusion criteria, and the use of examples. These variations reflected what aspects users prioritized when defining their criteria. To systematically analyze these differences, we generated a codebook and used it to categorize the observed criteria changes(see Table 4). We observed a range of specificity, with some users providing highly detailed and task-specific criteria, while others offered more general guidelines. In some cases, users added rules or additional prompting to guide the model. For example, some criteria included explicit if-then rules to distinguish between acceptable and unacceptable options. Additionally, some users incorporated examples within their criteria to illustrate preferred outcomes, effectively providing the model with concrete cases to guide its responses. This approach highlights the diverse ways in which participants interpreted and operationalized the constructs in question. Moreover, the number of items included in the scale for direct assessment also varied, where some users offered detailed multi-item scales, while others opted for simpler, binary options. .
5.2.1. Criteria Priority Differences within Tasks
Within each task, participants prioritized different aspects when defining and interpreting criteria, leading to varied approaches. Using inductive coding (Thomas, 2006), two authors generated a codebook of these priorities, as shown in Table 5.
@@ 5.2.1. Criteria Priority Differences within Tasks
In the Email Inclusivity task, participants prioritized cultural inclusivity, neutrality, and fairness. Some emphasized mentioning all cultures to avoid bias, ensuring holidays like Christmas, Hanukkah, Kwanzaa, and Diwali were included. Others focused on maintaining a neutral tone by avoiding specific mentions of culture, gender, race, or ethnicity. Additionally, some prioritized fairness, ensuring the email did not favor one culture over another. One participant intepreted inclusivity to mean financial inclusivity and designed the criteria around adherence to budget constraints, such as keeping gift exchange mentions within a \$40 limit.
@@
In the Summarization task, participants varied in their focus on factual accuracy, inclusion of key points, and conciseness. Some prioritized ensuring the summary accurately reflected the article, while others focused on covering all key points comprehensively. Conciseness was also important, with participants aiming for a summary around 20% of the original length. Grammatical correctness and the inclusion of crucial details, like dates or prices, were also emphasized. Some participants relied on the AI to decide which is the best summary by keeping their criteria general and asking for the the ”best” summary.
@@
For the Q&A Faithfulness task, participants focused on maintaining faithfulness to the source, adhering to policy, avoiding hallucinations, and ensuring correctness. They emphasized that responses should be strictly grounded in the reference document and follow company policy, with no deviation or fabrication. These differences within each task reflect the varied approaches participants took based on their interpretations and judgment of what was most important for the specific context.
@@
Additional Prompting The criteria includes instructions beyond criteria description Present Which summary offers the most clear and correct answer. As well as answering anything the customer did not know what to ask for in relation to their question. You must review each summary independently of the other summaries when making your judgement.
@@
5.3. RQ2: What strategies do users employ to refine their criteria for achieving human-AI alignment in LLM-as-a-judge evaluations?
To understand how people refine their criteria when using large language models (LLMs) as judges, we logged all metadata, including the criteria and AI evaluator used, for each evaluation across various tasks. Two authors coded the types of edits users made, identifying several ways participants modified their criteria to better align with task requirements or personal preferences. One common modification was making criteria more specific. For example, in the Email Inclusivity task, a participant refined the criterion from simply stating that ”The generated email should be inclusive” to specifying that it should use inclusive terms like ”holiday” and ”festive.” This added precision provided clearer guidance for evaluation. In contrast, some participants generalized their criteria to broaden their scope. In the same task, a criterion initially focused on specific groups—such as cultures and backgrounds—was simplified to a more general statement: ”The email is inclusive to all cultures and backgrounds.”
Participants also made adjustments by removing unnecessary scale items or adding more detailed options. For instance, in the Summarization task, an original scale with multiple levels (”Absolutely,” ”Somewhat,” ”No”) was simplified to a binary ”Yes” or ”No,” while in other cases, the scale was expanded to include detailed levels like ”Excellent,” ”Good,” ”Average,” and ”Poor.” In addition, some users specified existing scale items to clarify their meaning, such as adding definitions for ”Inclusive” and ”Not Inclusive” in the Email Inclusivity task. Conversely, others generalized scale items, broadening them to encompass a wider range of cultural practices. These results can be seen in Table 2.
Minor edits were also common, such as rephrasing criteria for clarity. For example, the question ”What is the best summary?” was revised to include clearer instructions: ”Which of the following summaries best describe the article? Review each independently of the others.” Some participants added exclusion criteria to prevent irrelevant elements from being considered. One participant, for instance, added a criterion stating that ”The summary should be accurate and concise. Has no fake data generated outside of reference,” clearly outlining what should be excluded. These refinements illustrate the range of strategies participants used to tailor their criteria, from increasing specificity to allowing greater flexibility. Figure 8 shows the sequential changes in evaluation criteria by task type and user.
@@
5.4. RQ3: How do task-related factors and judge strategy impact how practitioners refine criteria?
@@ 5.4.2. What are the differences in human-AI alignment?
We examined the the extent to which users agreed with the AI model’s assessments across different conditions. Upon completion of each task, users indicated agreement with the model’s judgement. We conducted a repeated measures ANOVA, focusing on the judge strategies (pairwise vs. rubric) and the tasks (Q&A Faithfulness, Email Inclusivity, Summarization). The analysis revealed a marginally significant main effect for task type, $F(2,28)=2.73,p=0.08$, indicating that there was a difference between alignment achieved between the tasks. Tukey post hoc analysis further confirmed this finding, showing a marginally significant difference in alignment between the the Q&A task and the Email Generation Task (mean difference = -0.633, p = 0.08). The number of agreements can be seen in Figure 10.
5.5. RQ4: How Do Task-Related Factors and Judge Strategy Impact User Perceptions?
Figure 11. No significant differences in user trust in the AI evaluator were found between the tasks or the AI judge strategies.
Figure 12. Explanations were percieved as more helpful in the rubric condition than in the pairwise condition (p¡0.05).
Figure 13. No significant differences in perceived positional bias helpfulness were found across tasks or AI judge strategies.
@@ 5.5. RQ4: How Do Task-Related Factors and Judge Strategy Impact User Perceptions?
We aimed to investigate whether task-related factors and judge strategy impact user perceptions, specifically focusing on AI trust, self-satisfaction, perception of explanations, and cognitive load. For each of these measures, we conducted a repeated measures ANOVA, and we report the results below.
5.6. AI Trust
For trust in the AI evaluator, a repeated measures ANOVA showed no significant differences between tasks or AI strategies (Figure 14). Despite this, we identified reasons why participants either trusted or distrusted the AI evaluator.
5.6.1. Reasons for Trust
Participants who reported higher trust (above the median score of 4) highlighted several factors, such as the alignment between AI judgments and their criteria, prior positive experiences with specific models, and cohesive explanations. Some examples of feedback include:
@@ 5.6.4. Perception of Positional Bias
Of the 196 total evaluations run by the 15 users, positional bias appeared in 35% of those evaluations. We conducted an ANOVA to determine whether the type of AI Judge strategy (pairwise vs. rubric) and task type (Q&A, email generation, summarization) affected participants’ perception of how helpful indication of positional bias was in completing the tasks. The analysis revealed no significant main effects, indicating that neither the type of task nor the strategy used by the judge had a significant impact on participants’ perceptions of the helpfulness of positional bias (see Figure 14). This lack of effect could potentially be explained by the fact that positional bias was only present in 35% percent of evaluations. An independent samples t-test was conducted to compare the perceived helpfulness when participants results included positional bias and when they did not. The results showed a significant difference, t(28) = 2.98, p = 0.003, showing that when positional bias was present, participants reported it being more helpful than it did not appear in the results. The primary reason for positional bias not being helpful was that it did not appear in the results.
@@ 5.6.5. Cognitive Load
5.7. RQ5: What do practitioners prefer: Direct Assessment or Pairwise?
Participants were asked to reflect on whether they preferred direct assessment via creating rubrics and whether they preferred pairwise, or whether it was contextual. Participant responses ranged with almost half (7 participants) reporting that they preferred direct assessment, and the remaining being split between pairwise and saying it depends/its contextual. The overview of preference reasons can be seen in Table 16. Below we report why participants preferred one judge strategy over another.
Preference Code Definition of the Code
Direct Clarity and Control Clearer and easier to understand, offering more control over the evaluation process and outcomes.
Detailed Evaluation Allows individual evaluation of responses, making it more useful for tasks requiring thorough review.
Pairwise Flexibility and Nuance Provides flexibility and nuance in evaluating abstract or subjective tasks.
Ease of Criteria Formulation Easier to create evaluation criteria, especially when rubrics are harder to define.
Neither Combination Approach A combination of both methods works better for complex tasks, complementing each other through classification (rubric) and ranking (pairwise).
Task-Specific Use Cases Direct Assessment is suited for classification or compliance, while Pairwise is better for identifying preferences or the best result.
Figure 15. Codes and Reasons for Assessment Preferences
Figure 16. Assessment Preferences
@@ 6.1.1. Mitigating Over-Specificity in AI-Assisted Evaluation: Balancing Task Context and Generalization
When analyzing how participants adjusted their criteria to align with the AI’s judgments, we observed a consistent tendency to make the criteria increasingly specific to the task at hand. This resulted in overly narrow criteria that worked well for evaluating a single article’s summary but failed to generalize across multiple summaries from different articles. Other AI-assisted evaluation systems (Kim et al., 2023b; Shankar et al., 2024) do not necessarily prioritize sampling diverse task contexts and outputs, especially when the evaluation heavily depends on prompts. This limits the user’s exposure to varied outputs, increasing the risk of overspecifying criteria to a particular task. To address this issue, it is critical to expose users to diverse task contexts. For instance, in the case of article summarization, this would mean providing users with varied articles and summaries to encourage the development of criteria that can be applied more broadly. A designimprovement could involve enabling users to upload large datasets, with the system recommending or allowing the selection of diverse samples to help develop more robust criteria.
Conversely, we found that a subset of users kept constructs very general. For example they would rely on the AI evaluator to interpret what it means to be “inclusive” or what it means to be grounded in the document. This lack of specificity can also lead to subpar results. Striking a balance between overly specific and overly vague criteria is essential. Encouraging users to refine their criteria while keeping them general enough for broad application can improve quality. Paired with exposure to diverse outputs, this approach can create more effective evaluation processes that work well across different contexts.

### v2

@@ 6.1. Participant Demographics
We recruited 15 participants at our company who had prior experience with model evaluation. The job titles are listed in Table 2. Their prior experience included subjective and manual evaluation, use case specific evaluation, evaluation framework and metrics.
@@
6.2. RQ1: How do participants approach the development and refinement of evaluation criteria when using EvalAssist, and what strategies, characteristics, and evaluative dimensions emerge in their criteria creation process?
In our first research question, we are interested in exploring the diversity of approaches leveraged by participants throughout the criteria development process. Because iteration is a key affordance of EvalAssist, we examined the strategies participants employed within tasks to refine their criteria (Table 3). Beyond the changes made during each evaluation within a task, we also investigated the characteristics of the criteria when participants reported satisfaction with their criteria at the completion of an evaluation task (Table 4). Lastly, we explored how participants prioritized certain evaluative dimensions in their criteria (e.g., faithfulness versus conciseness) and how they translated task instructions into actionable criteria for the AI evaluator (Table 5). To generate these insights, we employed inductive coding (Thomas, 2006). Two coauthors independently extracted themes from the data, focusing on the characteristics of the criteria generated for each task. They then met to compare and discuss their initial codes to identify areas of agreement and disagreement. Discrepancies were resolved through discussion, and additional rounds of refinement were conducted to ensure that the codes accurately captured the nuances of the data. This process culminated in the development of a consensus codebook, which was iteratively refined and systematically applied to the entire dataset to ensure reliability and consistency.
6.2.1. Strategies for Refining Criteria
Table 3 summarizes strategies used to refine criteria based on the coded meta-data from all logged evaluations. This data included the criteria and AI evaluators applied across various tasks. In the direct assessment context, the most common refinement strategy was to make definitions more specific (37%), followed by changes to the model (18%). For example, in the Email Inclusivity task, a participant refined the criterion ”The generated email should be inclusive” by specifying that it should use terms like ”holiday” and ”festive” instead of potentially non-inclusive language. This added precision offered clearer guidance for evaluation. Later, in Section 8.1.1, we explore these trends in detail, with a particular focus on how specificity influences criteria creation. Figure 9 depicts the sequential progression of changes. Figure 8 provides an example of how criteria evolved over time.
In the direct assessment context, participants also made refinements by simplifying or elaborating on scale items. For instance, in the Summarization task, a multi-level scale (”Absolutely,” ”Somewhat,” ”No”) was simplified to a binary scale (”Yes” or ”No”) for greater clarity. Conversely, some scales were expanded to include additional levels such as ”Excellent,” ”Good,” ”Average,” and ”Poor.” Additionally, participants clarified existing scale items by adding definitions. For example, in the Email Inclusivity task, they explicitly defined terms like ”Inclusive” and ”Not Inclusive.” In contrast, some users generalized scale items, broadening them to account for a wider range of cultural practices.
One common modification was making criteria more specific. In contrast, some participants generalized their criteria to broaden their scope. In the same task, a criterion initially focused on specific groups, such as cultures and backgrounds, was simplified to a more general statement: ”The email is inclusive to all cultures and backgrounds.” Minor edits were also common, such as rephrasing criteria for clarity. For example, the question ”What is the best summary?” was revised to include clearer instructions: ”Which of the following summaries best describe the article? Review each independently of the others.” Some participants added exclusion criteria to prevent irrelevant elements from being considered. One participant, for instance, added a criterion stating that ”The summary should be accurate and concise. Has no fake data generated outside of reference,” clearly outlining what should be excluded. These refinements illustrate the range of strategies participants used to tailor their criteria, from increasing specificity to allowing greater flexibility. Figure 9 shows the sequential changes in evaluation criteria by task type and user.
6.2.2. Characteristics of Finalized Criteria
@@
The final criteria that users created varied widely across tasks, not only in terms of specificity, but also in the presence of instructions or rules, the number of items in the scale (for the direct assessment condition), the inclusion of exclusion criteria, and the use of examples. These variations reflected what aspects users prioritized when defining their criteria (see Table 4). We observed a range of specificity, with some users providing highly detailed and task-specific criteria, while others offered more general guidelines. In some cases, users added rules or additional prompting to guide the model. For example, some criteria included explicit if-then rules to distinguish between acceptable and unacceptable options. This may be explained by users’ need for transparency, specifically, their inability to see exactly how the AI evaluator is being prompted, leading them to add their own additional prompts. For example, the inclusion of, “You must review each summary independently of other summaries when making your judgment.” in the criteria suggests that the user believed the summaries were not being evaluated independently. We discuss providing transparency to users in Section 8.1.2 based on these findings.
Additionally, some users incorporated examples within their criteria to illustrate preferred outcomes, effectively providing the model with concrete cases to guide its responses. The lack of transparency here fails to clearly communicate to the user that the model may not be incorporating such examples during the evaluation phase. This approach and others highlights the diverse ways in which participants interpreted and operationalized the constructs in question.
6.2.3. Prioritization of Evaluative Dimensions
Within each task, participants generated diverse criteria for each task reflecting varied prioritization of evaluative dimensions (see Table 5). Our tasks were purposefully diverse so that we can explore the varied prioritization across tasks.
@@ 6.2.3. Prioritization of Evaluative Dimensions
In the Email Inclusivity task, participants prioritized cultural inclusivity, neutrality, and fairness. Some emphasized mentioning all cultures to avoid bias, ensuring holidays like Christmas, Hanukkah, Kwanzaa, and Diwali were included. This approach reflected a belief in the importance of visibility for diverse groups. Others focused on maintaining a neutral tone by avoiding specific mentions of culture, gender, race, or ethnicity. Their goal was to maintain a universal tone that could resonate with a broader audience without alienation. Additionally, some prioritized fairness, ensuring the email did not favor one culture over another. One participant intepreted inclusivity to mean financial inclusivity and designed the criteria around adherence to budget constraints, such as keeping gift exchange mentions within a \$40 limit. By designing criteria that specified gift exchange limits (e.g., within a \$40 budget), this participant highlighted an often-overlooked dimension of inclusivity: financial fairness.
@@
In the Summarization task, participants varied in their focus on factual accuracy, inclusion of key points, and conciseness. Some prioritized ensuring the summary accurately reflected the article, while others focused on covering all key points comprehensively. Conciseness was also important, with participants aiming for a summary around 20% of the original length. Beyond these priorities, some participants leaned on grammatical correctness and the inclusion of specific details, such as dates or prices, to enhance clarity and utility. Some participants relied on the AI to decide which is the best summary by keeping their criteria general and asking for the the “best” summary. This strategy reflected trust in the system’s evaluative judgment.
@@
For the Q&A Faithfulness task, participants focused on maintaining faithfulness to the source, adhering to policy, avoiding hallucinations, and ensuring correctness. They emphasized that responses should be strictly grounded in the reference document and follow company policy, with no deviation or fabrication.
These varied approaches across tasks demonstrate the diversity in participant priorities and interpretations. Each decision reflected a combination of task-specific goals, individual values, and context-driven judgments, revealing the complexity of aligning evaluation criteria with multifaceted objectives. This richness in participant strategies highlights the importance of flexibility and transparency in criteria development to accommodate diverse perspectives.
Type of Change Original Criteria Changed Criteria
Scale: Absolutely, Somewhat, No Scale: Yes, No
Scale: Yes, No Scale: Excellent, Good, Average, Poor
Maybe: Not sure. Maybe: Not sure.
Minor Edit What is the best summary? Which is the best summary?
Table 3. Types of changes made across every evaluation.
@@
Additional Prompting The criteria includes instructions beyond criteria description Present Which summary offers the most clear and correct answer. As well as answering anything the customer did not know what to ask for in relation to their question. You must review each summary independently of the other summaries when making your judgment.
@@
6.3. RQ2: How do task-related factors and judge strategy impact how practitioners refine criteria?
@@ 6.3.2. What are the differences in human-AI alignment?
We examined the the extent to which users agreed with the AI model’s assessments across different conditions. Upon completion of each task, users indicated agreement with the model’s judgment. We conducted a repeated measures ANOVA, focusing on the judge strategies (pairwise vs. rubric) and the tasks (Q&A Faithfulness, Email Inclusivity, Summarization). The analysis revealed a marginally significant main effect for task type, $F(2,28)=2.73,p=0.08$, indicating that there was a difference between alignment achieved between the tasks. Tukey post hoc analysis further confirmed this finding, showing a marginally significant difference in alignment between the the Q&A task and the Email Generation Task (mean difference = -0.633, p = 0.08). The number of agreements can be seen in Figure 11.
6.4. RQ3: How Do Task-Related Factors and Judge Strategy Impact User Perceptions?
@@
Figure 13. Explanations were perceived as more helpful in the rubric condition than in the pairwise condition (p¡0.05).
We investigated whether task-related factors and judge strategy impact user perceptions, specifically focusing on AI trust, perception of explanations, and cognitive load. For each of these measures, we conducted a repeated measures ANOVA, and we report the results below.
6.4.1. AI Trust
For trust in the AI evaluator, a repeated measures ANOVA showed no significant differences between tasks or AI strategies. Despite this, we identified reasons why participants either trusted or distrusted the AI evaluator. Participants who reported higher trust (above the median score of 4) highlighted several factors, such as the alignment between AI judgments and their criteria, prior positive experiences with specific models, and cohesive explanations. Some examples of feedback include:
@@ 6.4.4. Perception of Positional Bias
Of the total evaluations run by the 15 users, positional bias appeared in 35% of those evaluations. We conducted an ANOVA to determine whether the type of AI Judge strategy (pairwise vs. rubric) and task type (Q&A, email generation, summarization) affected participants’ perception of how helpful indication of positional bias was in completing the tasks. The analysis revealed no significant main effects, indicating that neither the type of task nor the strategy used by the judge had a significant impact on participants’ perceptions of the helpfulness of positional bias. This lack of effect could potentially be explained by the fact that positional bias was only present in 35% percent of evaluations. An independent samples t-test was conducted to compare the perceived helpfulness when participants results included positional bias and when they did not. The results showed a significant difference, t(28) = 2.98, p = 0.003, showing that when positional bias was present, participants reported it being more helpful than it did not appear in the results. The primary reason for positional bias not being helpful was that it did not appear in the results.
@@ 6.4.5. Cognitive Load
6.5. RQ4: What do practitioners prefer: Direct Assessment or Pairwise Comparison?
Participants were asked to reflect on their preferred judgment strategy, whether direct assessment, pairwise comparison, or a context-dependent approach, and to explain the reasons behind their preference. Participant responses ranged with almost half (7 participants) reporting that they preferred direct assessment, and the remaining being split between pairwise and saying it depends/its contextual. The overview of preference reasons can be seen in Table 6. Below we report why participants preferred one judge strategy over another.
@@ 6.5.3. Use of AI Judge Strategy is Contextual
7. Limitations
While this study offers insights into how practitioners refine evaluation criteria and interact with AI-assisted judgment strategies, there are a few limitations. First, participants were recruited from a single organization. Although they had relevant domain expertise, this recruitment context may limit the generalizability of our findings. Future work should expand to more diverse populations, including external practitioners with varied evaluation needs and workflows. Additionally, even though EvalAssist was designed with scalability in mind through features such as task abstraction and flexible input configurations, its capabilities were only explored in small-scale, interactive scenarios in this study. Future work should test criteria generalization across larger datasets, and varied prompts. Third, the study relied primarily on participants’ self-reported perceptions of trust, cognitive load, and explanation quality. While subjective in nature, these responses offer valuable insight into users’ real-time judgments and experiences during evaluation. Explanations were also more accessible in the Direct Assessment condition than in the Pairwise condition, which may have influenced trust and perception. Future work can examine how explanation visibility shapes user trust and evaluation behavior.
@@ 8.1. Variability in Subjective Criteria: Defining Stakeholder Needs in AI-assisted Evaluation
The tasks assigned to users (email generation, summarization preference, and Q&A faithfulness), differed in their levels of specificity and ambiguity. However, we observed that regardless of how specific the instructions were, people consistently interpreted the task in significantly different ways. Ultimately, it is the human who defines the evaluation criteria and communicates these expectations to the AI-evaluator. These findings highlight the importance of careful deliberation and decision-making with all stakeholders and impacted parties when defining evaluation criteria. For example, stakeholders must explicitly decide what terms like ”inclusive” or ”a good summary” mean before relying on an LLM judge. Adopting deliberative processes like participatory design with diverse stakeholders, including end-users, can contribute to ensure multiple perspectives are considered. These steps ensure that the resulting criteria is aligned with the needs of all parties involved. Additionally, differences in alignment were observed across task types, suggesting that some tasks may naturally achieve better alignment than others. This underscores the need for machine learning practitioners to work with all relevant stakeholders to agree on clear, precise directions tailored to the specific task to ensure consistent and effective evaluations.
@@ 8.1.1. Mitigating Over-Specificity in AI-Assisted Evaluation: Balancing Task Context and Generalization
When analyzing how participants adjusted their criteria to align with the AI’s judgments, we observed a consistent tendency to make the criteria increasingly specific to the task at hand. This resulted in overly narrow criteria that worked well for evaluating a single article’s summary but failed to generalize across multiple summaries from different articles. Other AI-assisted evaluation systems (Kim et al., 2024a; Shankar et al., 2024) do not necessarily prioritize sampling diverse task contexts and outputs, especially when the evaluation heavily depends on prompts. This limits the user’s exposure to varied outputs, increasing the risk of overspecifying criteria to a particular task. To address this issue, it is critical to expose users to diverse task contexts. For instance, in the case of article summarization, this would mean providing users with varied articles and summaries to encourage the development of criteria that can be applied more broadly. A design improvement could involve enabling ML practitioners to upload large datasets, with the system recommending or allowing the selection of diverse samples to help develop more robust criteria.
Conversely, we found that a subset of users kept constructs very general. For example they would rely on the AI evaluator to interpret what it means to be “inclusive” or what it means to be grounded in the document. This lack of specificity can also lead to subpar results. Striking a balance between overly specific and overly vague criteria is essential. Encouraging users to refine their criteria while keeping them general enough for broad application can improve quality. Paired with exposure to diverse outputs, this approach can create more effective evaluation processes that work well across different contexts. Some participants over-specified their criteria by embedding ground-truth expectations, while others generalized too broadly. As previously discussed, the nature of the representative data influenced how participants refined their criteria. These observations show the importance of selecting diverse and representative examples, including multiple instances per task, to support effective and generalizable criteria refinement. This finding directly informs our design recommendations for improving LLM-as-a-judge systems.
Preference Code Definition
Direct Assessment Clarity and Control Offers a clearer, more understandable evaluation process with greater user control.
Detailed Evaluation Enables response-by-response evaluation, which is helpful for tasks requiring thorough review.
Pairwise Comparison Flexibility and Nuance Supports nuanced evaluation of abstract or subjective tasks.
Ease of Criteria Formulation Simplifies the creation of evaluation criteria when rubrics are difficult to define.
Neither (Contextual) Combination Approach A combined use of both methods is effective for complex tasks, balancing classification and ranking needs.
Task-Specific Fit Direct Assessment is suited for compliance/classification, while Pairwise is better for preference-based selection.
Table 6. Codes and rationales for users’ preferred evaluation strategies.